OpenStudy (anonymous):

one of sin x, cos x and tan x is given. Find the other two if it lies in the specified interval: tanx= -(3/4), x in [pi/2 , pi] The answer is: sin x= -(3/5), cos x = (4/5) Can someone help me solve this problem?? I don't understand it.

OpenStudy (anonymous):

first draw a picture in a graph of a triangle like so

OpenStudy (anonymous):

then find the length of the hypotenuse which is 5

OpenStudy (anonymous):

using that way you can solve out sine and cosine which are the answers you've already gave

OpenStudy (anonymous):

btw the signs of sine or cosine can be switched. so either sine = 3/5 and cosine = -4/5 or sine = -3/5 and cosine = 4/5 are correct

OpenStudy (anonymous):

if your teacher goes from the answers from a calculator though, you can only draw negative tan triangles in the fourth quadrant because that's how a calculator works. You could still argue the other answer is gave and be given credit for it though if you explain with the diagram above.

OpenStudy (anonymous):

Ok, yes!! I forgot! I was making it more complicated than it needed to be! Thanks!!